Bob Agrajag

Item Card  Freely Transeferable

0101 A Mop

A nameplate on the handle reads “Bob Agrajag”.

PROP

Hitchhiker’s Guide: B Ark  IN: Spring 1998

Item Card  Freely Transeferable

1234 A Golgafrinchan ID Card for “Bob Agrajag”

Opens type C locks.

Hitchhiker’s Guide: B Ark  IN: Spring 1998
Item Card  Freely Transeferable

5544 A bucket.

Holds 2 gallons of water  1 dot bulky
Hitchhiker’s Guide: B Ark  IN: Spring 1998